# Open dataset for benchmarking scaling laws of high-energy laser atmospheric propagation

### 14 Apr 2026 v1

@@ 4.1 Model-level Validation
Figure 3: Resultant beam intensity distribution in steady-state thermal blooming simulation for a 200-kW laser beam with 500 mm aperture and 5 km focus distance. Left column: intensity spots computed by the proposed method; right column: results reproduced from [14] for comparison.
@@
For thermal blooming validation, we simulate a steady-state propagation scenario with a transmitted power of 200 kW, an aperture diameter of 500 mm, and a focus distance of 5 km. The atmospheric absorption coefficient $\alpha_{{\textrm{abs}}}$} is set to $5\times{10^{-6}}$ m-1 and $1\times{10^{-5}}$ m-1, while $v_{{\textrm{wind}}}$} is set to 5 m/s and 10 m/s. The resulting beam intensity profiles are shown in Figure 3. A comparison with reference [14] simulations under identical conditions shows good agreement in both the beam shape and peak intensity. Figure 3 clearly illustrates the progressive distortion of the beam due to thermal blooming effects. In the baseline case (Figure 3a, e ), where both absorption and wind velocity are zero, the beam retains its near-diffraction-limited focus. As thermal blooming intensifies—driven by increased absorption coefficients and wind velocities—the beam becomes increasingly asymmetric, stretched in the wind direction, and exhibits reduced peak intensity. Figure 3b, f show that a moderate absorption with a wind velocity of 10 m/s leads to lateral beam shifting and intensity spreading. This trend becomes more pronounced in Figure 3c, d and g, h, where the higher absorption and smaller wind speed induces severe thermal lensing and results in increasing intensity profiles. Overall, our results (left column) are in good agreement with those of [14] (right column), both in structure and intensity scale, confirming the validity of the proposed simulation approach.
@@
Here, $a_{Y}$} is the 63.2% encircled energy radius at the target plane, and $F_{\text{a}}$ is the truncation ratio, defined as the ratio of the aperture diameter to the beam waist radius, $D/{a_{0}}$. For each aperture diameter $D$, we selected corresponding $a_{0}$} values to yield ${F_{a}}=D/{a_{0}}=$ 2.0, 2.8, 4.0, 5.0, 5.7 and 6.0, respectively. These $F_{\text{a}}$ values of were deliberately selected to match those used in Ref. [4], in order to facilitate direct comparison of accuracy. As shown in Figure 4, the fitted values of the scaling parameter $c_{3}$} under different truncation conditions are in close agreement with the reference results reported in [4].
@@ 5 Usage Notes
All data generated and described in this study are publicly available in Zenodo under DOI 10.5281/zenodo.19157773. The repository contains the complete simulation dataset, including the case-level tabular records, metadata files, variable dictionary, and the associated array files for long-exposure irradiance distributions. The dataset is distributed under the Creative Commons Attribution 4.0 International license. The simulation and data-conversion codes used in this study are available from the corresponding author upon reasonable request.

### 17 Apr 2026 v2

@@ 4.1 Model-level Validation
Figure 3: Resultant beam intensity distribution in steady-state thermal blooming simulation for a 200-kW laser beam with 500 mm aperture and 5 km focus distance. Left column: intensity spots computed by the proposed method; right column: results reproduced from Schleijpen[14] for comparison.
@@
For thermal blooming validation, we simulate a steady-state propagation scenario with a transmitted power of 200 kW, an aperture diameter of 500 mm, and a focus distance of 5 km. The atmospheric absorption coefficient $\alpha_{{\textrm{abs}}}$} is set to $5\times{10^{-6}}$ m-1 and $1\times{10^{-5}}$ m-1, while $v_{{\textrm{wind}}}$} is set to 5 m/s and 10 m/s. The resulting beam intensity profiles are shown in Figure 3. A comparison with reference [14] simulations under identical conditions shows good agreement in both the beam shape and peak intensity. Figure 3 clearly illustrates the progressive distortion of the beam due to thermal blooming effects. In the baseline case (Figure 3a, e ), where both absorption and wind velocity are zero, the beam retains its near-diffraction-limited focus. As thermal blooming intensifies—driven by increased absorption coefficients and wind velocities—the beam becomes increasingly asymmetric, stretched in the wind direction, and exhibits reduced peak intensity. Figure 3b, f show that a moderate absorption with a wind velocity of 10 m/s leads to lateral beam shifting and intensity spreading. This trend becomes more pronounced in Figure 3c, d and g, h, where the higher absorption and smaller wind speed induces severe thermal lensing and results in increasing intensity profiles. Overall, our results (left column) are in good agreement with those of Schleijpen[14] (right column), both in structure and intensity scale, confirming the validity of the proposed simulation approach.
@@
Here, $a_{Y}$} is the 63.2% encircled energy radius at the target plane, and $F_{\text{a}}$ is the truncation ratio, defined as the ratio of the aperture diameter to the beam waist radius, $D/{a_{0}}$. For each aperture diameter $D$, we selected corresponding $a_{0}$} values to yield ${F_{a}}=D/{a_{0}}=$ 2.0, 2.8, 4.0, 5.0, 5.7 and 6.0, respectively. These $F_{\text{a}}$ values of were deliberately selected to match those used in Chen’s work [4], in order to facilitate direct comparison of accuracy. As shown in Figure 4, the fitted values of the scaling parameter $c_{3}$} under different truncation conditions are in close agreement with the reference results reported in Chen’s work[4].
@@ 5 Usage Notes
All data generated and described in this study are publicly available in Zenodo under DOI 10.5281/zenodo.19157773. The repository contains the complete simulation dataset, including the case-level tabular records, metadata files, variable dictionary, and the associated array files for long-exposure irradiance distributions. The dataset is distributed under the Creative Commons Attribution 4.0 International license. The simulation and data-conversion codes used in this study cannot be shared at this time because the code also forms part of an ongoing paper.